A car accelerates uniformly from rest to a speed of 36 km/hr in 5 seconds. The distance covered by the car in the first 10 seconds is:


1
Expert's answer
2021-06-07T09:34:57-0400

"v=at\to a=v/t=10/5=2\ (m/s^2)"


"s=at^2/2=2\cdot10^2/2=100\ (m)" - if the car moves with acceleration within "10 \ 
s"


"s=at^2/2+vt=2\cdot5^2/2+10\cdot5=75\ (m)" - if the car moves with acceleration for "5 \ 
s" , and


evenly for the next "5 \ s" .